OpenStudy (anonymous):

Use a proportion to answer the question in terms of y. What number is 15% of y???????

OpenStudy (tkhunny):

\(15\)%\(=\dfrac{15}{100} = \dfrac{Fred}{y}\) Solve for Fred?

OpenStudy (anonymous):

i need a number not a name

OpenStudy (tkhunny):

That is not correct. You need a name. You have no value for 'y'. Why do you think Fred can be replaced by a number? Read the problem statement carefully: "...in terms of y."

OpenStudy (anonymous):

then what number is 15% of y?

OpenStudy (tkhunny):

Use that proportion I gave you. Solve for "Fred". Trust me on this.

OpenStudy (anonymous):

We have only done this kind of math with numbers never with a name. So i don't know how I am suppose to do the math

OpenStudy (tkhunny):

Well, without a value for y, there will be no numbers, today. Q: What difference does it make if it's a name or a number? A: None, whatsoever. Just solve for Fred. Maybe you don't like Fred? He's a little too glib? Okay, fine... \(15\)%\(=\dfrac{15}{100} = \dfrac{x}{y}\) Solve for x. Just solve for x. Don't worry about the meaning of the result, just yet. Just solve for x.